– TUESDAY

Registration 9.00-10.00 Plenary talks:ROOM D (SISSA Basement)

10 : 00 − 10 : 45 Saito 11 : 15 − 11 : 45 Gibbons 11 : 45 − 12 : 30 Elbau

Lunch Break

14 : 30 − 15 : 15 Santini 15 : 30 − 16 : 15 Konopelchenko

Session 1: ROOM D

16 : 45 − 17 : 15 Mañas 17 : 15 − 17 : 45 Moro 18 : 00 − 18 : 30 Carlet

Session 2: ROOM B

16 : 45 − 17 : 15 Takemura 17 : 15 − 17 : 45 Spiridonov 18 : 00 − 18 : 30 Novokshenov

WEDNESDAY

Plenary talks:

9 : 30 − 10 : 15 Bobenko 10 : 45 − 11 : 30 Cattaneo 11 : 45 − 12 : 30 Kuijlaars 14 : 30 − 15 : 15 Hertling 15 : 30 − 16 : 15 Veselov

Session A:

16 : 45 − 17 : 15 Marshall 17 : 15 − 17 : 45 Daems 18 : 00 − 18 : 30 Duits

Session B:

16 : 45 − 17 : 15 Dunajski 17 : 15 − 17 : 45 Hitzazis 18 : 00 − 18 : 30 Arnlind

THURSDAY

Plenary talks:

9 : 30 − 10 : 15 Zuber 10 : 45 − 11 : 30 Smirnov 11 : 45 − 12 : 30 Kazakov 14 : 30 − 15 : 15 Shramchenko 15 : 30 − 16 : 15 Constantin 16 : 45 − 17 : 30 Hoppe

FRIDAY

Plenary talks: Lecture Hall (SISSA main entrance)

9 : 30 − 10 : 15 Fokas 10 : 45 − 11 : 30 van Moerbeke 11 : 45 − 12 : 30 Tracy 14 : 30 − 15 : 15 Novikov 15 : 30 − 16 : 15 Crowdy

Session 1: MAIN Lecture Hall

16 : 45 − 17 : 15 Ragnisco 17 : 15 − 17 : 45 Lorenzoni

Session B: ROOM D

16 : 45 − 17 : 15 Klein 17 : 15 − 17 : 45 Kamvissis

MAIN Lecture HALL SATURDAY

Plenary talks:

9 : 30 − 10 : 15 Martinez-Alonso 10 : 45 − 11 : 30 Ferapontov 11 : 45 − 12 : 30 Samaj 14 : 30 − 15 : 15 Suris 15 : 30 − 16 : 15 Mason

Session 1:

MAIN Lecture Hall

16 : 45 − 17 : 15 Khimshiashvili 17 : 15 − 17 : 45 Vanlessen 17 : 45 − 18.15 Garcia Estevez

Session 2:

Room D

16 : 45 − 17 : 15 Strachan 17 : 15 − 17 : 45 Hone 17 : 45 − 18 : 15 Hallnäss